# A switching identity for cable-graph loop soups and Gaussian free fields

### v3

@@ Abstract.
We derive a “switching identity” that can be stated for critical Brownian loop-soups or for the Gaussian free field on a cable graph: It basically says that at the level of cluster configurations and at the more general level of the occupation time fields, conditioning two points on the cable-graph to belong to the same cluster of Brownian loops (or equivalently to the same sign-cluster of the GFF) amounts to adding a random odd number of independent Brownian excursions between these points to an otherwise unconditioned configuration. This explicit simple description of the conditional law of the clusters when a connection occurs has various direct consequences, in particular about the large scale behaviour of these sign-clusters on infinite graphs.
@@ 1. Introduction
Let us make the following first comments:
@@
    Introducing random walk representation of fields, lattice models and exploiting some of their combinatorial features is an idea that can be traced back (at least) to the seminal works of Symanzik or Brydges, Fröhlich and Spencer [58, 7], see [35] for a recent relevant account. The present switching property can be reminiscent of the switching property for the random current representation of the Ising model from Griffith, Hurst and Sherman [24] later developed and used very fruitfully by Aizenman and others, starting with [3] (see the “the random current revolution” section in the review [19]) where it is explained how Ising correlations (and more) could therefore be expressed in terms of probabilities involving multiple independent currents. We stress however that – as is apparent from our GFF cable-graph statements here, where the events that one conditions on are connection events on the cable graph – our results are very much “cable-graph ones” (or more generally “continuum ones”, see Section 6.3), and do not seem to have an as simple counterpart on discrete graphs, despite of the relations such as the ones studied in [40], so that they can not be accessed so naturally via the toolboxes that have been developed for discrete models.
@@
    One first main step in the proof will be to see that when conditioning on $\Lambda(x)$ and $\Lambda(y)$ only (and not on the fact that $x\leftrightarrow y$), then the decomposition of the Brownian loops that hit both $x$ and $y$ in the loop-soup into excursions away from $\{x,y\}$ combined with the rewiring property will yield a similar decomposition as in Theorem 2, but where the final item is replaced by a Poisson point process of excursions with intensity $ab\nu_{x,y,G}$ that is conditioned to be even (instead of odd). Note that with this description, if one then additionally conditions on $x\leftrightarrow y$, the four inputs become very much correlated in the case where there is no excursion joining $x$ and $y$ (and when $x$ and $y$ are far away from each other, this will be the dominating event) since the event $x\leftrightarrow y$ will typically be created by an excursion away from $x$ that touches either an excursion away from $y$ or a chain of loops that then touches an excursion away from $y$. This description with even number of excursions does therefore not provide that much insight about connectivity events.
    It will turn out that the probability that if we now consider instead the same four independent items but with an unconditioned (i.e., it can be even or odd) Poisson number of excursions joining $x$ and $y$, then the probability of having an odd number of excursions is in fact be identical to the probability of having an even number of excursions and a connection between $x$ and $y$ (if one adds the first three items). The theorem therefore shows that in this setup, there exists a (measure-preserving) bijection that preserves the occupation times between the configurations with an odd number of excursions and the configurations with an even number of excursions that do create a connection between $x$ and $y$. This bijection is the “switching” in the name that we gave to this property.
@@
    While the existence of this bijection between configurations of (even number of excursions from $x$ to $y$ + loop-soup) conditioned to connect $x$ and $y$ and configurations of (odd number of excursions from $x$ to $y$ + loop-soup) follows from the theorem, neither the statement nor the proof that we will give does provide an explicit construction or even insight into what such a bijection might look like. Some thought does actually lead to the idea that such an explicit bijection cannot be obtained by finitely many reconnections/rewiring of loops/excursions, due to parity issues. The arguments and results of [51, 32] also provide food for thought in this direction. They indeed suggest that if the switching corresponds to some sort of rewiring on the loops and excursions, then some exceptional points that were not part of a chain of loops in the first instance will become part of an excursion (note that if these points are sufficiently sparse, then this won’t affect the total occupation time).
    In Section 5, we will describe heuristically and explicitly one of the possible bijections. This hopefully enlightens further what is going on here (even if the coupling is by no means trivial). This will build on the limit of reverse-vertex reinforced jump process that has been studied by Lupu, Tarrès and Sabot [39]. This approach and its consequences will be detailed in subsequent work [65, 42].
@@
    One can a posteriori detect aspects of this general switching property in a number of recent works on the GFF, especially those with the loup-soup perspective. This includes in the continuous and conformally invariant two-dimensional setting the papers [51, 32] that we will further comment on, as well as [25] [Remark 1.12 in that paper can be viewed as the scaling limit i.e., as the continuum version of Theorem 1 in this two-dimensional setting]. In the cable-graph setting, as we shall discuss in Section 3.3, an identity in law about bridges of three-dimensional Bessel processes by Pitman and Yor from the 1982 paper [47] can actually be interpreted as the switching property for the cable graph consisting of just one edge. Another relevant example is the paper [1] (see in particular Theorem 1.4 there that one could then use by adding an artificial additional edge bewtween two points to a cable graph) that also builds on the considerations [62, 40] (as part of the present paper does) that describes the number of excursions along one edge in the cable graph, where the conditioned Poisson random variables appear.
@@
    Both the GFF and its square (via the loop-soup representation) do satisfy a spatial Markov property. The switching property in some sense allows to reconcile the apparent parity-type contradiction that the combination of these two Markov properties do seem to create. Of course, one can also view it the other way around (and this is one way the proofs will go), namely that the switching property is a consequence of the combination of these two Markov properties.
@@
The paper is structured as follows: In Section 2, we will prove Theorem 2 directly via loop-soup considerations. In Section 3, we describe another approach/interpretation to this proof in the spirit of Dynkin’s isomorphism theorem. In Section 4, we will describe the aforementioned direct consequences of the switching property (the constructions and descriptions of various versions of the infinite incipient clusters and statements about the combination of loop-soup with interlacements). In Sections 5 and 6, we will briefly outline some forthcoming work and results (respectively on the explicit bijection where the even/odd bijection, and on some further consequences about the asymptotic behaviour of large clusters, for instance on how to extract information about multiple-point functions).
@@
Let us conclude this introduction with some informal comments about why such a striking general result has not been uncovered before. Indeed, once one guesses that the statement holds, it is not really difficult to prove (and to come up with different proofs)! And as already mentioned, one can a posteriori see it “not far from the surface” at various places in the literature. The main reason is arguably simply that the bijection between configurations with even and with odd number of excursions is not trivial (which provides an excuse to the community – or at least to me – to why it was not so easy to guess the result), especially when one looks at it from the lens of Dynkin’s isomorphism theorem without its loop-soup interpretation. Also, as we have already mentioned, it is not an observation that comes from discrete models – ideas from the continuum are needed. We note that it is not the first time that work on the (conformally invariant) two-dimensional continuous settings did lead mathematicians on the trail of general basic results about the GFF and loop-soups on cable graphs: The loop-soup itself was first introduced in the continuum setting in [31] in relation to the scaling limit of loop-erased random walk and the conformal restriction properties developed in [29], and its relation to the continuum two-dimensional GFF (via the combination of [54, 44] and [55]) was derived before Le Jan [33, 34] pointed out the direct relation between the square of the discrete GFF and the discrete loop-soups in general transient graphs or the direct basic relation with the loops erased in Wilson’s algorithm was worked out (see e.g. [28, 66]). This also holds in the present case: In [32], we were zooming in on some decompositions of two-dimensional loop-soup clusters given part of their boundaries (following earlier considerations in [51, 49, 50]), where some quizzing parity issues naturally popped up. A similar-looking switching identity appeared in [32] [in the latter section of that paper, we derive an identity in the (more involved) continuous two-dimensional case and for rectangular domains with some very special boundary conditions, so the points $x$ and $y$ in some sense correspond to two opposite vertical sides of the rectangle and the excursions from $x$ to $y$ do correspond to horizontal crossings]. This in turn led to the trail of the type of bijection that we will describe in Section 5 and also to the realisation that this type of switching result was actually working already at the simpler level of cable-graph GFFs and not only for some specific values of the values $a$ and $b$. So, again a somewhat convoluted route – with a special role played by the considerations, arguments and questions raised in [51, 32] coming from the continuum two-dimensional world.
2. Proof of the switching property
@@
We now turn to the proof of Theorem 2. Let us first write some words about the way in which we normalize the various objects involved: The GFF is defined via the Green’s function of the cable-graph Brownian motion, which is the expected local time at $y$ cumulated by a Brownian motion started at $x$ until the possibly infinite time at which it exits the graph. The normalization of the Brownian loop measure is then the one such that the occupation time of a Poisson point process with that intensity is exactly distributed as the square of the GFF (this corresponds to $c=1$ in the normalization/notation of [31] inspired by the notion of central charge for two dimensional models, or $\alpha=1/2$ in the papers by Le Jan and Lupu). The excursion measure $\nu_{x,x}=\nu_{x,x,G}$ away from a point $x$ in $G$ is chosen in such a way that $\nu_{x}(\ell_{e}(z))=\phi_{x}(z)^{2}$ where $\ell_{e}$ is the local time of the excursion $e$ at $z$ and $\phi_{x}(z)$ is the harmonic function in $G\setminus\{x\}$ with boundary values $1$ at $x$ and 00 at all other boundary points (and at infinity). The excursion measure $\nu_{x,y}$ between two points can then be constructed as the measure on paths joining $x$ to $y$, obtained by first restricting the measure $\nu_{x,x}$ to those paths that hit $y$, and then keeping only the part of this path up until the first time it hits $y$.
@@
The first step of the proof will be to derive the following fact that gives the complete picture for Description (2):
@@
Remark 5.
One could (as in [32], where the analog of a special instance of this result in the two-dimensional continuous setting is derived) prove a more general statement with $n$ points $x_{1},\ldots,x_{n}$: The conditioning would then be that if $N_{i,j}=N_{j,i}$ is the number of excursions joining $x_{i}$ to $x_{j}$ in $G\setminus\{x_{1},\dots,x_{n}\}$, then $(N_{i,j})_{1\leq i<j\leq n}$ has the law of independent Poisson random variables with respective means $a_{i}a_{j}$ times the mass of the corresponding excursion measure, where this collection is conditioned by the event that for all $i$, $\sum_{j\not=i}N_{i,j}$ is even. But for the purpose of the present paper, the case $n=2$ is all what we need.
Before giving the proof of the lemma, let us first show how to use it in order to deduce the switching result. The basic idea is to notice that, when one conditions on $x\leftrightarrow y$, then the law of the configuration described in (1) i.e., via an unconditioned Poisson random variable $N$, is the same as the law of the configuration obtained via (2), i.e., when the Poisson random variable is conditioned to be even. As a consequence, the same is true when one conditions on the remaining event $\{x\leftrightarrow y\}\setminus\{N\hbox{ is even}\}$, i.e., when $N$ is conditioned to be odd.
@@
We now turn our attention to the proof of Lemma 4: We will derive it as a consequence of results for loop-soups of a simple continuous-time Markov chain $X$ with two states $x$ and $y$ (and a cemetery boundary state $\partial$) defined as follows: When at $x$, the chain jumps to $y$ with rate $\alpha>0$ and it dies with rate $u_{x}$. When at $y$, the chain jumps to $x$ with the same rate $\alpha$ and it dies with rate $u_{y}$. Indeed, when following the trace of the cable-graph Brownian motion on the set $\{x,y\}$, which is all one is interested when counting its excursions between $x$ and $y$, one has exactly a Markov chain of this type (the time of the latter is the local time accumulated by the former at $x$ or $y$).
@@
Our proof will be based on the ideas of Proposition 2.46 from [66] (see also [28] for analogous statements), which gives the explicit expression for the number of jumps on edges for discrete loop-soups (with discrete time). We each (large) constant $K$, we define the discrete-time Markov chain $X^{K}$ on the same set $\{x,y,\partial\}$ with jump probabilities $p(x,y)=p(y,x)=\alpha/K$, $p(x,x)=1-(u_{x}+\alpha)/K$, $p(x,\partial)=u_{x}/K$, $p(y,\partial)=u_{y}/K$, $p(y,y)=1-(u_{y}+\alpha)/K$ (the point $\partial$ is an absorbing state). In the limit when $K\to\infty$, the number of steps of this discrete chain $X^{K}$ divided by $K$ then corresponds to the time for the continuous-time Markov chain $X$.
@@
The paper [1] by Elie Aidékon contains some closely related considerations, see in particular his Theorem 1.4 where parity shows up – it does not address the switching property in terms of conditioning on long connections, but one a posteriori detect the switching property hiding near-by – one can for instance try to see what happens when one adds an additional edge to his setting. As we shall see, he also looked into the direction that we will describe in Section 5.

### v4

@@ Abstract.
We derive a “switching identity” that can be stated for critical Brownian loop-soups or for the Gaussian free field on a cable graph: It basically says that at the level of cluster configurations and at the more general level of the occupation time fields, conditioning two points on the cable-graph to belong to the same cluster of Brownian loops (or equivalently to the same sign-cluster of the GFF) amounts to adding a random odd number of independent Brownian excursions between these points to an otherwise unconditioned configuration. Such explicit simple descriptions of the conditional law of the clusters when a connection occurs have various direct consequences, in particular about the large scale behaviour of these sign-clusters on infinite graphs.
@@ 1. Introduction
We will also discuss switching results along closed loops, that are more reminiscent of the switching identities for the random current representation of the Ising model as initiated by Aizenman [3]. Here we come back to the definition of the cable-graph based on a discrete graph (with sites $S$ and edges $E$) and view $G$ as the union of the segments corresponding to the edges of $E$. We can note that each edge/segment will be crossed only finitely many times by loops in the Brownian loop-soup ${\mathcal{L}}$. The loops are not oriented, but we can make sense of the parity $P(e)\in{\mathbb{Z}}/(2{\mathbb{Z}})$ of the total number of crossings of an edge by ${\mathcal{L}}$ (i.e., the sum over all loops of ${\mathcal{L}}$ of their number of crossings of the edge); as opposed to the actual number of crossings of the edge, this parity remains unchanged if one glues two Brownian loops into one, or splits one Brownian loop into two loops. Let $E_{\Lambda}$ denote the set of edges along which $\Lambda$ remains strictly positive, and let $E(P)$ denote the set of edges for which $P$ is equal to $1$ (so this means that the number of crossings is odd). Clearly, $E(P)\subset E_{\Lambda}$ (because if $\Lambda$ hits 00 on the edge, there can not be any crossing), and $(S,E(P))$ is an even subgraph of $(S,E_{\Lambda})$ (i.e., for each site, the number of edges of $E(P)$ adjacent is even – this is just because the crossings are made by loops).
Proposition 4.
Consider a Brownian loop-soup on the cable graph $G$ and let $\Lambda$ be its occupation time field. Conditionally on $\Lambda$, the graph $(S,E(P))$ is a uniform random even subgraph of $(S,E_{\Lambda})$.
[When a graph is infinite but has only finite connected components, we say that a uniform even random subgraph is obtained by choosing independently a uniform random even subgraph for each component.] This proposition is reminiscent of ideas from [44], but one important feature to stress here is that this result is valid conditionally on all of $\Lambda$ (not just on $E_{\Lambda}$). Since the conditional law depends on $E_{\Lambda}$ only, this means that conditionally on $E_{\Lambda}$, the field $\Lambda$ and $P$ are independent. So, one can first observe the random set $E_{\Lambda}$, then choose independently the random even subgraph, and independently of the random even subgraph, the values of $\Lambda$ at the sites and then all the conditionally independent values of $\Lambda$ on each edge. As we shall see, this corresponds to the fact that if one conditions on $\Lambda$, there is a “measure-preserving switch” of the parity of all the edges along any given cycle of edges contained in $E_{\Lambda}$ (and this switch does not change $\Lambda$).
Let us describe one interesting concrete consequence of this proposition here: Consider the case of a cable graph that is contained in ${\mathbb{Z}}^{d}$ (and that we view as embedded in ${\mathbb{R}}^{d}$), and consider the $(d-2)$-dimensional space $\Delta:=\{1/2\}\times\{1/2\}\times{\mathbb{R}}^{\{d-2\}}$ (so when $d=2$, this is the point $(1/2,1/2)$, when $d=3$, it is a line and so on). Then, each oriented loop in the cable graph has an index around $\Delta$ (which is the index around $(1/2,1/2,0,\ldots,0)$ of its projection on the plane ${\mathbb{Z}}^{2}\times\{0\}^{d-2}$). Whether this index is even or odd does not change when one changes the orientation of the loop. So, one can define unambiguously the total parity index $P_{\Delta}(C,{\mathcal{L}})$ of the collection of loops in ${\mathcal{L}}$ that are part of a cluster $C$ of loops as the sum (in ${\mathbb{Z}}/2{\mathbb{Z}}$) of the parity of the indexes of each of the Brownian loops in the cluster (which is finite as there are only finitely many loops that visit more than one site in $C$). Note that this quantity is again invariant under rewiring of the unoriented Brownian loops (i.e., it does not change when one concatenates or splits finitely many Brownian loops). Note finally that if $C$ does not contain any self-avoiding cycle of odd index around $\Delta$, then $P_{\Delta}(C,{\mathcal{L}})$ is necessarily equal to 00.
Corollary 5.
Conditionally on $\Lambda$ (which contains the information defining all the loop-clusters), the collection $(P_{\Delta}(C,{\mathcal{L}}))$ for the loop soup clusters that do contain cycles with an odd index around $\Delta$ form an i.i.d. family of Bernoulli-(1/2) random variables (i.e., one tosses one fair coin for each of these clusters to choose its loop-soup index parity).
This result will work also if one replaces $\Delta$ by a simple compact $(d-2)$-dimensional manifold that does intersect the cable-graph. One somewhat surprising consequence is that regardless of its shape, a loop-soup cluster that wraps around $\Delta$ (i.e., contains a cycle with non-zero index around $\Delta$) will have a conditional probability at least $1/2$ to actually contain a Brownian loop that wraps around $\Delta$ (indeed, if $P_{\Delta}(C,{\mathcal{L}})=1$, then at least one of the Brownian loops in $C$ must has a non-zero index). This will allow to prove Lupu’s “intensity doubling conjecture” formulated in [41] (i.e., to prove that it holds) about loop-soups in high dimensions. See Section 4 (and [46]) for more on all these aspects.
Let us make the following first comments related to these statements:
@@
[…]
@@
    Introducing random walk representation of fields, lattice models and exploiting some of their combinatorial features is an idea that can be traced back (at least) to the seminal works of Symanzik or Brydges, Fröhlich and Spencer [62, 8], see [38] for a recent relevant account. The present switching property can be reminiscent of the switching property for the random current representation of the Ising model from Griffith, Hurst and Sherman [25] later developed and used very fruitfully by Aizenman and others, starting with [3] (see the “the random current revolution” section in the review [20]) where it is explained how Ising correlations (and more) could therefore be expressed in terms of probabilities involving multiple independent currents. We stress however that – as is apparent from our GFF cable-graph statements here, where the events that one conditions on are connection events on the cable graph – our results are very much “cable-graph ones” and do not seem to have an as simple counterpart for discrete loop-soups. On the other hand (see Section 4), it will be possible to relate aspects of our switching identities to those appearing in the random current representations of the Ising model (as the results of [44] also suggest).
@@
    We will provide several proofs. Let us say a few words about one rather “self-contained” proof to explain where the switching is coming from: One first main step will be to see that when conditioning on $\Lambda(x)$ and $\Lambda(y)$ only (and not on the fact that $x\leftrightarrow y$), then the decomposition of the Brownian loops that hit both $x$ and $y$ in the loop-soup into excursions away from $\{x,y\}$ combined with the rewiring property will yield a similar decomposition as in Theorem 2, but where the final item is replaced by a Poisson point process of excursions with intensity $ab\nu_{x,y,G}$ that is conditioned to be even (instead of odd). Note that with this description, if one then additionally conditions on $x\leftrightarrow y$, the four inputs become very much correlated in the case where there is no excursion joining $x$ and $y$ (and when $x$ and $y$ are far away from each other, this will be the dominating event) since the event $x\leftrightarrow y$ will typically be created by an excursion away from $x$ that touches either an excursion away from $y$ or a chain of loops that then touches an excursion away from $y$. This description with even number of excursions does therefore not provide that much insight about connectivity events.
    However, one can also look at the same conditional law for $\Lambda$ when one adds an extra ghost edge between $x$ and $y$ along which the GFF does not hit 00. The Markov property of the GFF shows that one obtains the same conditional distribution for $\Lambda$ but with an extra weighting of $2$ for the configurations for which $x$ is not connected to $y$ in $G$ (as one does not need to toss a coin so that they have the same sign). So, this reweighted law can then be described by the same four independent items but with an unconditioned (i.e., it can be even or odd) Poisson number of excursions joining $x$ and $y$ in $G$. Comparing this with the previous description then allows to conclude that the probability of having an odd number of excursions is in fact be identical to the probability of having an even number of excursions and a connection between $x$ and $y$ (if one adds the first three items), and lead to the theorem.
@@
    The theorem shows the existence of a (measure-preserving) bijection that preserves the occupation times between the configurations with an odd number of excursions and the configurations with an even number of excursions that do create a connection between $x$ and $y$. This bijection is the “switching” in the name that we gave to this property. However, neither the statement nor the idea of proof that we just gave do provide an insight into what such a bijection might look like. Some thought does actually lead to the idea that such an explicit bijection cannot be obtained by finitely many reconnections/rewiring of loops/excursions, due to parity issues. The arguments and results of [55, 35] also provide food for thought in this direction. They indeed suggest that if the switching corresponds to some sort of rewiring on the loops and excursions, then some exceptional points that were not part of a chain of loops in the first instance will become part of an excursion (note that if these points are sufficiently sparse, then this won’t affect the total occupation time).
    In Section 4, we will discuss two aspects of these bijections: One at the level of edges (leaving a black box about what happens within each edge) that is closely related to the usual switching for random currents, and one at the level of the individual Brownian loops that hopefully enlightens what is going on here (even if the coupling is by no means trivial). This will build on the limit of reverse-vertex reinforced jump process that has been studied by Lupu, Tarrès and Sabot [43]. This approach and its consequences will be detailed in subsequent work [69, 46].
@@
    Both the GFF and its square (via the loop-soup representation) do satisfy a spatial Markov property. The switching property in some sense allows to reconcile the apparent parity-type contradiction that the combination of these two Markov properties do seem to create. Of course, one can also view it the other way around (and this is basically the way the proof outlined above goes), namely that the switching property is a consequence of the combination of these two Markov properties. In some sense, all three proofs can be viewed as variations around this theme.
@@
The paper is structured as follows: In Section 2, we will describe a proof Theorem 2 that is fairly direct and self-contained, and builds mostly on loop-soup considerations and the rewiring property in particular. In Section 3, we describe another approach/interpretation to this proof in the spirit of Dynkin’s isomorphism theorem, so this involves more some Gaussian random variables. Section 4 provides yet another approach that highlights the link with the random current representation of the Ising model (and leads to Proposition 4 and to the proof of Lupu’s intensity doubling conjecture). We also informally discuss in that section how to get an explicit bijections at the level of the Brownian loops. In Section 5, we then describe the aforementioned direct consequences of the switching property (the constructions and descriptions of various versions of the infinite incipient clusters and statements about the combination of loop-soup with interlacements). In Section 6, we briefly outline some forthcoming work and results (respectively on the explicit bijection at the level of Brownian loops, and on some further consequences about the asymptotic behaviour of large clusters, for instance on how to extract information about multiple-point functions).
Let us conclude this introduction with the following important remark: The proof (or the proofs, since we will provide several ones) of these switching identities are not technically difficult, and the underlying ideas bear many similarities with many of the papers in the area. In fact, once one realizes that the statements do hold, it is actually easy to work out proofs, building on various items in the existing literature – and each of those acquainted with them will surely quickly be able to provide their own prefered approach (this includes Yves Le Jan or Lorca Heeney who independently sent me some compact versions after the posting of the first version of this paper, based on the results of [44, 38]). One can in fact detect aspects of this general switching property in a number of recent works on the GFF, especially those with the loup-soup perspective. This includes in the continuous and conformally invariant two-dimensional setting the papers [55, 35] that we will further comment on, as well as [27] [Remark 1.12 in that paper can be viewed as the scaling limit i.e., as the continuum version of Theorem 1 in this two-dimensional setting] or Lupu’s paper on the twisted GFF [40]. In the cable-graph setting, as we shall discuss in Section 3.3, an identity in law about bridges of three-dimensional Bessel processes by Pitman and Yor from the 1982 paper [51] can actually be interpreted as the switching property for the cable graph consisting of just one edge. Another relevant example is the paper [1] (see in particular Theorem 1.4 there that one could then use by adding an artificial additional edge bewtween two points to a cable graph) that also builds on the considerations [66, 44] (as part of the present paper does) that describes the number of excursions along one edge in the cable graph, where the conditioned Poisson random variables appear. So, one possible short proof is to build on [66, 44] to quickly deduce the switching for one edge, and then the arguments in Section 4 that are more “discrete switching” arguments in the spirit of [4] and random current representations, to deduce the result in general cable graphs.
@@
One can therefore wonder why such a striking general result that has many nice consequences has not been uncovered before. The main reason is arguably simply that the explicit bijection between configurations with even and with odd number of excursions at the level of the Brownian paths is not trivial (which provides an excuse to the community – or at least to me – to why it was not so easy to guess the result), especially when one looks at it from the lens of Dynkin’s isomorphism theorem without its loop-soup interpretation. Also, as we have already mentioned, it is not an observation that comes from discrete models – ideas from the continuum (at least on an interval) are needed to get the ball rolling. We note that it is not the first time that work on the (conformally invariant) two-dimensional continuous settings did lead mathematicians on the trail of general basic results about the GFF and loop-soups on cable graphs: The loop-soup itself was first introduced in the continuum setting in [34] in relation to the scaling limit of loop-erased random walk and the conformal restriction properties developed in [32], and its relation to the continuum two-dimensional GFF (via the combination of [58, 48] and [59]) was derived before Le Jan [36, 37] pointed out the direct relation between the square of the discrete GFF and the discrete loop-soups in general transient graphs or the direct basic relation with the loops erased in Wilson’s algorithm was worked out (see e.g. [31, 70]). This also holds in the present case: In [35], we were zooming in on some decompositions of two-dimensional loop-soup clusters given part of their boundaries (following earlier considerations in [55, 53, 54]), where some quizzing parity issues naturally popped up. A similar-looking switching identity appeared in [35] [in the latter section of that paper, we derive an identity in the (more involved) continuous two-dimensional case and for rectangular domains with some very special boundary conditions, so the points $x$ and $y$ in some sense correspond to two opposite vertical sides of the rectangle and the excursions from $x$ to $y$ do correspond to horizontal crossings]. This in turn led to the trail of the type of bijection that we will describe in Section 4.4 and also to the realisation that this type of switching result was actually working already at the simpler level of cable-graph GFFs and not only for some specific values of the values $a$ and $b$. So, again a somewhat convoluted route – with a special role played by the considerations, arguments and questions raised in [55, 35] coming from the continuum two-dimensional world.
2. A first proof of the switching property
@@
We now turn to a first proof of Theorem 2, which is the one that we outlined in the introduction. Let us first write some words about the way in which we normalize the various objects involved: The GFF is defined via the Green’s function of the cable-graph Brownian motion, which is the expected local time at $y$ cumulated by a Brownian motion started at $x$ until the possibly infinite time at which it exits the graph. The normalization of the Brownian loop measure is then the one such that the occupation time of a Poisson point process with that intensity is exactly distributed as the square of the GFF (this corresponds to $c=1$ in the normalization/notation of [34] inspired by the notion of central charge for two dimensional models, or $\alpha=1/2$ in the papers by Le Jan and Lupu). The excursion measure $\nu_{x,x}=\nu_{x,x,G}$ away from a point $x$ in $G$ is chosen in such a way that $\nu_{x}(\ell_{e}(z))=\phi_{x}(z)^{2}$ where $\ell_{e}$ is the local time of the excursion $e$ at $z$ and $\phi_{x}(z)$ is the harmonic function in $G\setminus\{x\}$ with boundary values $1$ at $x$ and 00 at all other boundary points (and at infinity). The excursion measure $\nu_{x,y}$ between two points can then be constructed as the measure on paths joining $x$ to $y$, obtained by first restricting the measure $\nu_{x,x}$ to those paths that hit $y$, and then keeping only the part of this path up until the first time it hits $y$.
@@
The first step of the proof is the following fact that gives the complete picture for Description (2):
@@
Actually, a more general statement with $n$ points $x_{1},\ldots,x_{n}$ holds: The conditioning would then be that if $N_{i,j}=N_{j,i}$ is the number of excursions joining $x_{i}$ to $x_{j}$ in $G\setminus\{x_{1},\dots,x_{n}\}$, then $(N_{i,j})_{1\leq i<j\leq n}$ has the law of independent Poisson random variables with respective means $a_{i}a_{j}$ times the mass of the corresponding excursion measure, where this collection is conditioned by the event that for all $i$, $\sum_{j\not=i}N_{i,j}$ is even. In fact, this general result (and therefore also this parity lemma) is almost exactly Proposition 7 in [66]. The one slight difference in the formulation there is also that it describes only the conditional law of the excursions from $x$ to $y$ (or from any $x_{i}$ to any $x_{j}$ for $i\not=j$), and does not say anything about the excursions from one of the points to itself. It is however not difficult (for instance building on the discrete setting) to see that these point processes of excursions will be independent (and of course the same for the loops that do not intersect the marked points, since they are anyway independent from $(\Lambda(x_{i}))_{i\leq n}$). Since the details of the proof there were “left as a simple exercise” (i.e., the part that derives it from its analog for discrete time loop-soups), we will provide some more details about this proof.
But before that, let us first show how to use it in order to deduce the switching result. The basic idea is to notice that, when one conditions on $x\leftrightarrow y$, then the law of the configuration described in (1) i.e., via an unconditioned Poisson random variable $N$, is the same as the law of the configuration obtained via (2), i.e., when the Poisson random variable is conditioned to be even. As a consequence, the same is true when one conditions on the remaining event $\{x\leftrightarrow y\}\setminus\{N\hbox{ is even}\}$, i.e., when $N$ is conditioned to be odd.
@@
We now turn our attention to the proof of Lemma 6: We will detail the proof outlined in [66]. A first step is to view this as a result for loop-soups of a simple continuous-time Markov chain $X$ with two states $x$ and $y$ (and a cemetery boundary state $\partial$) defined as follows: When at $x$, the chain jumps to $y$ with rate $\alpha>0$ and it dies with rate $u_{x}$. When at $y$, the chain jumps to $x$ with the same rate $\alpha$ and it dies with rate $u_{y}$. Indeed, when following the trace of the cable-graph Brownian motion on the set $\{x,y\}$, which is all one is interested when counting its excursions between $x$ and $y$, one has exactly a Markov chain of this type (the time of the latter is the local time accumulated by the former at $x$ or $y$).
@@
Our proof will be based on the ideas of Proposition 2.46 from [70] or [66] (see also [31] for analogous statements), which gives the explicit expression for the number of jumps on edges for discrete loop-soups (with discrete time). We each (large) constant $K$, we define the discrete-time Markov chain $X^{K}$ on the same set $\{x,y,\partial\}$ with jump probabilities $p(x,y)=p(y,x)=\alpha/K$, $p(x,x)=1-(u_{x}+\alpha)/K$, $p(x,\partial)=u_{x}/K$, $p(y,\partial)=u_{y}/K$, $p(y,y)=1-(u_{y}+\alpha)/K$ (the point $\partial$ is an absorbing state). In the limit when $K\to\infty$, the number of steps of this discrete chain $X^{K}$ divided by $K$ then corresponds to the time for the continuous-time Markov chain $X$.
@@
The paper [1] by Elie Aïdékon contains some closely related considerations, see in particular his Theorem 1.4 where parity shows up – it does not address the switching property in terms of conditioning on long connections, but one a posteriori detect the switching property hiding near-by – one can for instance try to see what happens when one adds an additional edge to his setting. As we shall see, he also looked into the direction that we will describe in Section 4.4.
Remark 11.
We can note that with this approach, the proof for a general $G$ does not differ much from the case where the only connection between $x$ and $y$ goes along a single edge of the graph – replacing the entire graph by a single edge with the same “effective resistance” (which corresponds to ideas developed in [45]) works well in a bijective way, both in terms of excursions and loop-soups (and the parity of excursions between $x$ and $y$), the connection via $\Lambda$ or the GFF correlation functions.